Each of 39 of my friends has either a dog a cat or a rabbit each of 24 of them has a dog each of 17 gas a cat and each of 16 has a rabbit the number having both a dog and a cat is one more than the number having both a cat and a rabbit while 2of them have all the three how many of my friends have both a rabbit and cat


1
Expert's answer
2022-01-12T18:42:14-0500

Each of 39 of my friends has either a dog, a cat or a rabbit. Each of 24 of them has a dog, each of 17 has a cat and each of 16 has a rabbit. The number having both a dog and a cat is 1 more than the number having both a cat and a rabbit. There are 9 who have both a dog and a rabbit while 2 of them have all the three. How many of my friends have both a rabbit and a cat?

Let "x=" the number having both a cat and a rabbit. Then "x+1=" the number having both a dog and a cat.


"N(D\cup C\cup R)=N(D)+N(C)+N(R)"

"-N(D\cap C)-N(D\cap R)-N(C\cap R)"

"+N(D\cap C\cap R)"

"39=24+17+16-(x+1)-x-9+2"

"2x=10"

"x=5"

There are 5 who have both a rabbit and a cat.